### v1

# Distributional Random Forests for Complex Survey Designs on Reproducing Kernel Hilbert Spaces

Abstract
We study estimation of the conditional law $P(Y|X=\mathbf{x})$ and continuous functionals $\Psi(P(Y|X=\mathbf{x}))$ when $Y\in\mathcal{Y}$ takes values in a locally compact Polish space, $X\in\mathbb{R}^{p}$, and the observations arise from a complex survey design. We propose a survey-calibrated distributional random forest (SDRF) that incorporates complex-design features via a pseudo-population bootstrap, PSU-level honesty, and a Maximum Mean Discrepancy (MMD) split criterion computed from kernel mean embeddings of Hájek-type (design-weighted) node distributions. We provide a framework for analyzing forest-style estimators under survey designs; establish design consistency for the finite-population target and model consistency for the super-population target under explicit conditions on the design, kernel, resampling multipliers, and tree partitions. As far as we are aware, these are the first results on model-free estimation of conditional distributions under survey designs. Simulations under a stratified two-stage cluster design provide finite sample performance and demonstrate the statistical error price of ignoring the survey design. The broad applicability of SDRF is demonstrated using NHANES: We estimate the tolerance regions of the conditional joint distribution of two diabetes biomarkers, illustrating how distributional heterogeneity can support subgroup-specific risk profiling for diabetes mellitus in the U.S. population.
1 Introduction and Motivation
Experimental design is a key driver of successful statistical practice [97]. A rigorous study design protects and makes transparent the assumptions that link the data to scientific conclusions, thus improving the reliability and reproducibility of the empirical findings [71].
Complex survey design [42] methods provide a powerful methodology for population-representative inference in a wide range of settings, such as biomedical studies [59]. In the finite-population regime [25], a probability sample with known inclusion probabilities and carefully constructed weights allows design-unbiased (or asymptotically design-unbiased) estimation of population functionals. These estimators can exhibit satisfactory statistical properties, such as efficiency, provided that appropriate statistical calibration and inference procedures are used. This design-based framework is particularly compelling when paired with randomized interventions: probability surveys deliver external validity (representativeness of a well-defined target population), while randomized experiments deliver internal validity (causal identification under the design). Together, they support the policy-relevant generalization of causal effects [45].
Large, nationally representative health studies, such as the National Health and Nutrition Examination Survey (NHANES), which monitors health and nutrition in the U.S. population, are canonical examples of the impact of complex survey methodology in practice [68]. When design information is properly incorporated and analyzed with appropriate statistical methods, such data can be used to quantify disease burden, monitor disparities, and inform clinical and public health policy in a cost-effective way, all while ensuring valid population inference.
Gap and motivation: Machine learning and survey data.
Despite the central role of complex survey designs in nationally representative studies, predictive tools for complex surveys remain limited in two main respects. First, most existing methods target simple functionals such as conditional means or medians, rather than the full conditional distribution, and are typically based on linear models or classical nonparametric smoothers such as Nadaraya–Watson estimators and spline-based methods [61]. Second, available procedures often presuppose low-dimensional covariates and focus on smoothing over relatively simple or very smooth function classes. By contrast, the use of modern machine-learning methods that can capture complex structures more effectively in high-dimensional settings such as neural networks [5, 98] or tree-based ensembles [26]-—within the survey context remains relatively scarce [66, 67].
From an application’s perspective, many scientific questions demand distributional targets beyond means or medians [50], including multivariate quantiles, tolerance regions [34], tail probabilities, and other functionals of multivariate conditional laws. In NHANES, for example, multivariate distribution estimates can be used to define tolerance regions that detect outlying biomarker profiles given subject-specific characteristics. Treating complex-survey observations as i.i.d. data discards design information, degrading calibration and invalidating design-based uncertainty quantification, while properly incorporating weights and design features preserves both consistency and representativeness.
Modern machine-learning methods, such as neural networks and regression-tree ensembles, are effective in many high-dimensional regimes and can capture intricate statistical associations that are difficult to model with classical nonparametric techniques. However, they are typically formulated under i.i.d. sampling and, when naively applied to complex surveys, can distort node-wise conditional distributions, bias risk estimates, and compromise inference. A design-aware approach must propagate weights and clustering through splitting and aggregation, respect replicate-weight or bootstrap schemes that mimic the original selection, and produce distributional outputs for vector-valued $Y$, not merely conditional means. These requirements call for distributional learning procedures that integrate survey design into both the objective function and the algorithmic construction.
From a machine-learning perspective, integrating predictive algorithms with complex survey designs creates opportunities to increase the statistical efficiency of estimators, mitigate selection bias, reduce overfitting through design-guided regularization, improve subgroup fairness, and achieve proper predictive calibration in inference tasks in applied settings.
Our paper. We study nonparametric estimation of multivariate conditional distributions under complex survey designs. Let $(Y,X)$ denote a pair consisting of a vector-valued response $Y\in\mathcal{Y}$ and covariates $X\in\mathbb{R}^{p}$ observed through a probability sample drawn under a complex survey design. We introduce a design-aware distributional random forest estimator $\widehat{\mathbb{P}}_{Y\mid X=\mathbf{x}}$ that combines survey-weighted tree construction and aggregation with loss functions induced by mean kernel embeddings of probability measures in a separable reproducing kernel Hilbert space. This yields a single forest-based procedure that (i) targets the full conditional law $\mathbb{P}(Y\mid X=\mathbf{x})$ and, as plug-in estimators, a broad class of component-wise and joint functionals such as conditional quantiles, cross-moments, covariance matrices and tail probabilities, and (ii) remains consistent under the joint law induced by the super–population model and the survey design.
On the theoretical side, we formulate the estimation problem in terms of the maximum mean discrepancy (MMD) [37, 80, 89], a family of statistical distances that metrizes weak convergence in suitable classes of probability measures, and show that the forest weighing the survey achieves design and model consistency; that is, the MMD between $\widehat{\mathbb{P}}_{Y\mid X=\mathbf{x}}$ and $\mathbb{P}_{Y\mid X=\mathbf{x}}$ converges uniformly to zero over $x$ at rates governed by tree-partition and design complexity. We obtain a decomposition of the MMD risk into design, algorithmic, and stochastic terms, prove that each component vanishes under mild moment conditions on inclusion and replicate weights. Our analysis quantifies the impact of the number of trees $B$ and the bootstrap or replicate-weight draws $M$ on node-level predictions and on the variance of the ensemble, showing a reduction in variance of order $O(M^{-1/2})$ while preserving the consistency of the design. We emphasize that, even in the i.i.d. setting, the theory of classical random forests is far from trivial [9, 10] and has required substantial effort to develop since Breiman’s original proposal of random forests [17]. Moreover, many questions remain open, for instance regarding finite-sample guarantees [55] or extensions to dependent data [36].
We complement the theory with an application to NHANES, where we estimate the joint conditional distribution of two biomarkers used in diabetes diagnosis based on individual-specific characteristics, providing distributional estimates that can inform a range of decisions from a healthcare and public-health perspective.
Positioning within the literature.
Our work is based on two strands. On the survey side, foundational contributions make the sampling mechanism explicit through survey weights [96, 84, 40, 77, 30, 82, 81, 39], including the Horvitz–Thompson and Hájek estimators [44, 28], as well as design-aware procedures widely used, for instance, in model checking, hypothesis testing, and regression [51, 75, 61, 74, 70, 47]. On the machine-learning side, we exploit the statistical learning properties of random forests [17] to reduce variance estimation and even bias in many regimes via the bagging step [15, 55]. Random forests [17] and their distributional variants (through kernel mean embeddings and MMD) provide flexible nonparametric tools that achieve competitive performance in high-dimensional settings across many benchmarks [31] and can be used efficiently for two-sample testing, variable selection, and clustering, in addition to predictive tasks. However, existing random forest algorithms are largely developed for i.i.d. data   design-aware random forest methodology and theory for complex surveys remain underdeveloped.
From a regression-model perspective, existing survey methods focus primarily on parametric and semiparametric estimation using (pseudo) probability weights and estimating equations related to M- and Z-estimation [91, 69]. Classical nonparametric smoothers—including local polynomials [48], Nadaraya–Watson estimators [43], and spline-based additive models—have been adapted to complex surveys [14]. However, most of this literature focuses on continuous outcomes [83, 46] and conditional learning; comparatively less attention has been paid to multivariate outcomes and full conditional learning.
From a theoretical perspective, design-based central limit theorems are available for a variety of survey designs [6, 60, 78, 11]. Empirical-process formulations replace the uniform empirical measure with Horvitz–Thompson/Hájek-weighted measures; under suitable weight conditions, the resulting processes inherit many i.i.d. properties. Survey-aware resampling is well–developed [76]: bootstrap and replicate-weight methods with asymptotic guaranties approximate sampling distributions and can be adapted to distributional functionals. However, existing theoretical developments are not directly tailored to distributional random forests under complex designs, where the interaction between algorithmic randomness, survey weights, and distributional loss functions requires a novel theoretical survey framework. In this paper, we provide a design-aware distributional random forest for survey data, together with a corresponding theoretical framework for its design-based and model-based properties. When learning a conditional distribution, using design-aware algorithms is not only about potential efficiency gains; it is essential for statistical consistency, as we illustrate in Table S3.
The remainder of the paper is organized as follows. Section 2 formalizes the complex-survey setup and recalls background information on the mean kernel embeddings. Section 3 presents the algorithm and the main theoretical results. Section 4 reports simulations and comparisons with a distributional random forest that ignores the design. Section 5 analyzes NHANES. Section 6 discusses limitations and future work. All proofs of the theoretical results and a notation table (Table S3) are provided in the Supplementary Material.
2 Set-up and Background
Suppose that we observe a sample $D=\{(X_{i},Y_{i},w_{i})\}_{i=1}^{n_{s}}$ drawn from a finite population $D^{N}=\{(X_{i},Y_{i},Z_{i})\}_{i=1}^{N}$ of size $N$ under a sampling design $\mathbf{p}$ that may depend on auxiliary design variables $Z_{i}$, which we will formalize shortly. Here, $Y_{i}\in\mathcal{Y}$ is the outcome, $X_{i}\in\mathbb{R}^{p}$ are covariates and $w_{i}\in\mathbb{R}^{+}$ are survey weights (possibly incorporating nonresponse adjustment and calibration). Let $P^{N}_{Y\mid X=\mathbf{x}}$ denote the conditional distribution of a finite population induced by $D^{N}$. Let $P_{Y\mid X=\mathbf{x}}$ denote the super-population conditional law. Consider a fixed covariate value $\mathbf{x}\in\mathbb{R}^{p}$ and a measurable functional $\Psi:\mathcal{M}_{b}(\mathcal{Y})\to(\mathcal{V},d_{v})$ taking some probability measures on $\mathcal{Y}$ in some separable metric space $(\mathcal{V},d_{v})$, and it is continuous at the target. Our objective is to construct an estimator $\widehat{\Psi}_{N}(\mathbf{x})$ that is design-consistent for the finite-population target $\Psi(P^{N}_{Y\mid X=\mathbf{x}})$ and model-consistent for the super-population target $\Psi(P_{Y\mid X=\mathbf{x}})$. We will work on the joint probability space $(\Omega\times\mathcal{S}^{N},\,\mathcal{A}\otimes\sigma(\mathcal{S}^{N}),\,P_{\Omega\times\mathcal{S}^{N}})$. Specifically, denote $P_{\mathcal{S}^{N}|\omega}$ the measure in the design space, we want that for any $\varepsilon>0$,
$P_{S^{N}\mid\omega}\!\left(d_{v}\!\left(\widehat{\Psi}_{N}(\mathbf{x}),\Psi(P^{N}_{Y\mid X=\mathbf{x}})\right)>\varepsilon\right)\to 0\quad\text{$P_{\Omega}$--a.e.\,(design consistency), and }$
$P_{\Omega\times\mathcal{S}^{N}}\!\left(d_{v}\!\left(\widehat{\Psi}_{N}(\mathbf{x}),\Psi(P_{Y\mid X=\mathbf{x}})\right)>\varepsilon\right)\to 0\quad\text{ (model consistency). }$
2.1 Survey Design
We follow the finite population framework considered by [13, 41, 42]: Consider a sequence of finite populations $\{\mathcal{U}^{N}\}_{N\geq 1}$. Let $U^{N}:=\{1,\dots,N\}$, the index set for each $\mathcal{U}^{N}$ and $\mathcal{S}^{N}:=\{S:S\subset U^{N}\}$ the collection of subsets of $U^{N}$. Let $\sigma(\mathcal{S}^{N})$ denote the $\sigma$-algebra generated by $\mathcal{S}^{N}$. We view the finite population $\mathcal{U}^{N}$ as a realization of i.i.d. random vectors $\{(X_{i},Y_{i},Z_{i})\}_{i=1}^{N}$ in a common probability space $(\Omega,\mathcal{A},P)$, where $Z_{i}\in\mathcal{Z}$ denotes the auxiliary (design) information available in the frame. Let $D^{N}:=\{(X_{i},Y_{i})\}_{i=1}^{N}$ denote the finite-population data and write $Z^{N}:=(Z_{1},\dots,Z_{N})$. For any region ${A}(\mathbf{x})\subset\mathcal{X}$ indexed by some fixed $\mathbf{x}\in\mathcal{X}$, denote the empirical conditional measure induced by ${D}^{N}$ as
${P}^{N}_{Y|X\in A(\mathbf{x})}(\mathbf{x})=\frac{1}{N(\mathbf{x})}\sum_{i=1}^{N}I(X_{i}\in{A}(\mathbf{x}))\;\delta_{Y_{i}},\text{ where $\delta_{Y_{i}}$ the Dirac measure,}$
$N(\mathbf{x})=\sum_{i\in[N]}I(X_{i}\in A(\mathbf{x}))$. A sampling scheme can be formalized as a transition function $\bm{p}:\Omega\times\sigma(\mathcal{S}^{N})\mapsto[0,1]$. From $\bm{p}$, fix any $\omega\in\Omega$, we define a probability measure in the design space $(\mathcal{S}^{N},\sigma(\mathcal{S}^{N}))$ as $P_{\mathcal{S}^{N}}(A):=\sum_{S\in A}\bm{p}({Z}^{N}(\omega),S)$, $\forall A\in\sigma(\mathcal{S}^{N})$. We can then uniquely define a product probability measure in the product space $(\Omega\times\mathcal{S}^{N},\mathcal{A}\otimes\sigma(\mathcal{S}^{N}))$ as
$P_{\Omega\times\mathcal{S}^{N}}(E\times S):=\int_{E}\bm{p}(\omega,\{S\}){P}(d\omega),\hskip 8.5359pt\forall E\times\{S\}\text{ in simple rectangles in }\mathcal{A}\otimes\sigma(\mathcal{S}^{N}).$
Thus, a random sample $S\sim P_{\mathcal{S}^{N}}$ can be obtained by ${\xi}^{N}=(\xi_{1},\dots,\xi_{N})$, a random vector on $\mathcal{S}^{N}$, where $\xi_{i}:=\mathbbm{1}_{(i\in S)}\in\{0,1\}$. Define first-order inclusion probabilities $\pi_{i}(\omega):=\mathbb{E}_{\mathcal{S}}(\xi_{i}|Z(\omega))=\sum_{\{S:i\in S\}}\bm{p}(S,{Z}^{N}(\omega))$, and second-order inclusion probabilities $\pi_{ij}(\omega):=\mathbb{E}(\xi_{i}\xi_{j}|Z(\omega))$. Denote $n_{s}:=card(S)=\sum_{i=1}^{N}\xi_{i}$, the size of the random sample $S$ that is a random variable. We do not have access to the entire finite population of size $N$, but we work with a sample of size $n_{s}$. This leads to estimators that use weights to construct consistent estimators ([56]). For example, the Hájek (HJ) estimator for the true conditional probability ${P}^{N}_{Y|X\in A(\mathbf{x})}(\mathbf{x})$ in the design space is
$\widehat{P}^{N}_{HJ,Y|X\in A(\mathbf{x})}(\mathbf{x})=\frac{1}{\widehat{N}(\mathbf{x})}\sum_{i=1}^{N}\frac{\xi_{i}}{\pi_{i}}I(X_{i}\in A(\mathbf{x}))\;\delta_{Y_{i}},\quad\widehat{N}(\mathbf{x})=\sum_{i=1}^{N}\frac{\xi_{i}}{\pi_{i}}I(X_{i}\in A(\mathbf{x})).$
The SDRF estimator, which we denote as $\widehat{\Psi}_{N}(\mathbf{x})$, adopts the form Hájek, but with critical modifications from design-compatible resampling and an adaptive data-driven partition $L^{*}(\mathbf{x})$ in place of $A(\mathbf{x})$. We assume the following with respect to the design.
  • On study design
  • (D1)
    Conditionally non-informative design: $\xi^{N}\mathrel{\perp\!\!\!\perp}Y^{N}\mid(X^{N},Z^{N})$.
  • (D2)
    Stable sampling fraction: $n_{s}/N\rightarrow_{p_{d}}f$ for some $f\in(0,1]$, and $P_{\mathcal{S}^{N}|\omega}(n_{s,N}/N\leq\delta)\leq CN^{-\kappa}$, with some $\kappa>4$.
  • (D3)
    Random Sampling: There exist constants $0<\underline{\lambda}\leq\bar{\lambda}<\infty$ such that for any $N\geq 1$, $\sup_{i\in[N]}\pi_{i}\leq\bar{\lambda}n/N$, $\inf_{i\in[N]}\pi_{i}\geq\underline{\lambda}n/N$, where $n=\mathbb{E}_{P_{{\mathcal{S}_{N}}\mid\omega}}(n_{s}(\omega))$ the expected sample size.
  • (D4)
    Controlled second order: For any $N\geq 1$, $\max_{1\leq i\leq j\leq N}|\pi_{ij}-\pi_{i}\pi_{j}|\leq C_{1}/n_{s,N}$ a.s.
Remark.
Assumptions (D1)–(D4), or minor variants thereof, are standard regularity conditions in the survey-sampling literature. (D2)–(D4) imply the design-based law of large numbers and central limit theorem statements, summarized in Lemma 2.1. These probabilistic bounds will be invoked repeatedly to establish the consistency of $\widehat{\Psi}_{N}(\mathbf{x})$. (D1) allows us to separate the super–population randomness from the design randomness and to treat the design expectations conditionally on $(X^{N},Z^{N})$. (D2) imposes a stable sampling fraction and rules out designs with a decreasing sample size. (D3) assumes that the inclusion probabilities of first order are uniformly of order $n/N$ and bounded away from 0 and $1$. This prevents extreme survey weights. (D4) is common [12, 11, 19] and a mild second order condition that controls the dependence among the inclusion indicators $\{\xi_{i}\}_{i=1}^{N}$ to be of order $1/n_{s}$.
Lemma 2.1.
Under (D2)-(D4), we have that w.r.t the design law $P_{\mathcal{S}^{N}|\omega}$, $\frac{1}{N}\sum_{i=1}^{N}\!\left(\frac{\xi_{i}}{\pi_{i}}-1\right)\xrightarrow{p_{d}}0$, and $\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\!\left(\frac{\xi_{i}}{\pi_{i}}-1\right)=O_{p_{d}}(1).$
Simple random sampling without replacement (SRSWOR), Poisson (Bernoulli) sampling, SRSWOR, PPS with replacement with MOS bounded above and away from zero, and their stratified versions are all examples that satisfy (D1)–(D4). Two–stage designs that satisfy a counter-part of (D1)–(D4) at each of the two stages satisfy also (D1)–(D4) at the subject level (Lemma A.2, Supplementary Material).
2.2 Embedding of Measures
Kernel mean embeddings (KMEs) [38, 85] provide a flexible and conceptually simple way to represent probability distributions as single points in a reproducing kernel Hilbert space (RKHS) [1]. Since these representations are defined purely through expectations of the kernel evaluations, all operations on distributions are reduced to linear operations in the RKHS. As a result, questions about probability measures–such as whether two random variables share the same distribution–can be reformulated as questions about distances between vectors and efficiently evaluated via the kernel trick. The distance between two embedded distributions in the RKHS is called the maximum mean discrepancy (MMD). It is a fully nonparametric way to compare distributions, and for characteristic kernels it is equal to zero only when the two distributions are the same. This framework supports a wide range of kernel based methods for two-sample testing, independence testing, clustering, biclustering and related problems; see, e.g., [86, 87, 37, 80, 33, 67].
From the perspective of constructing random forests for distributional targets, incorporating KMEs into the algorithm offers several practical advantages: (i) It provides a distribution-level target, enabling random forests to discriminate conditional laws via an MMD-based splitting criterion. (ii) Empirical MMD can be computationally advantageous relative to alternative statistical distances, such as bounded Lipschitz or Wasserstein distances, for example, when $\mathcal{Y}=\mathbb{R}^{d}$ and $d$ is of moderate size [32]. (iii) From an implementation standpoint, using MMD is simple: it only requires changing how distances between random objects in the ambient space are computed (e.g., in a general Hilbert space or, more generally, in a metric space of negative type, as in the case of spherical data [62]). As a result, modifications to a standard random forest algorithm are minimal, even for outcomes in abstract spaces. (iv) Most importantly, because the KME is defined directly on probability measures, it applies equally to both finite-population and super-population conditional laws under standard regularity conditions, allowing us to analyze a broad class of sampling designs within a single metric framework.
We now introduce the corresponding concepts and definitions from a mathematical perspective.
Let $\mathbf{k}:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ be a bounded positive definite kernel with RKHS $(\mathcal{H},\langle\cdot,\cdot\rangle_{\mathcal{H}})$ and the associated canonical feature map $\Phi_{\mathbf{k}}(y):=\mathbf{k}(y,\cdot)\in\mathcal{H}$. For any $P\in\mathcal{M}_{b}(\mathcal{Y})$ in the space of finite signed Radon measures, define the kernel mean embedding as the mapping
$\mu_{\mathbf{k}}:\mathcal{M}_{b}(\mathcal{Y})\mapsto\mathcal{H},\hskip 5.69046pt{P}\mapsto\int_{\mathcal{Y}}\mathbf{k}(y,\cdot)d{P}(y)\in\mathcal{H}.$
so that by the reproducing property, for any $f\in\mathcal{H}$, $P\in\mathcal{M}_{b}(\mathcal{Y})$, $\mu_{\mathbf{k}}(P)=\int_{\mathcal{Y}}\Phi_{\mathbf{k}}(y)dP(y)=\mathbb{E}_{Y\sim P}[\Phi_{\mathbf{k}}(Y)]\in\mathcal{H},$ a mean element in the RKHS. This yields a natural distance in the RKHS as $d_{\mathbf{k}}(P,\;Q)=\|\mu_{\mathbf{k}}(P)-\mu_{\mathbf{k}}(Q)\|_{\mathcal{H}},\;P,\;Q\in\mathcal{M}_{b}(\mathcal{Y})$. We call $\mathbf{k}$ characteristic if $\mu_{\mathbf{k}}$ is injective. In that case, $d_{\mathbf{k}}(P,\;Q)=0$ if and only if $P=Q$, so $d_{\mathbf{k}}$ is a valid statistical distance for the unique identification of the probability measures.
For our consistency arguments, we require, in addition, that $d_{\mathbf{k}}$ metrizes weak convergence on $\mathcal{M}_{b}(\mathcal{Y})$. That is, for any sequence $P_{1},\dots\in\mathcal{M}_{b}(\mathcal{Y})$, $P_{N}\rightsquigarrow P\iff d_{\mathbf{k}}(P_{N},\;P)\to 0$, as $N\to\infty$. We call kernels that satisfy this property universal (initially proposed in [88]). This property concerns the approximation ability of the KME sample estimate to the true target as $N\to\infty$. We place the following assumptions:
  • On Space $\mathcal{Y}$ and Kernel $\mathbf{k}$
  • (S1)
    $\mathcal{Y}$ is a locally compact Polish space.
  • (K1)
    $\mathbf{k}$ is uniformly bounded, positive definite, and $C_{0}$-continuous, where $C_{0}$-continuous means $\mathbf{k}(\cdot,y)\in C_{0}(\mathcal{Y})$ for all $y\in\mathcal{Y}$, where $C_{0}(\mathcal{Y})$ denotes the Banach space of bounded continuous functions vanishing at infinity (Section 2 of [86]).
  • (K2)
    $\mathbf{k}$ is $c_{0}$-universal. Meaning that the RKHS $\mathcal{H}$ induced by $\bm{k}$ satisfying (K1) is dense in $C_{0}(\mathcal{Y})$ defined above w.r.t. the uniform norm (Definition 2, Theorem 1 in [20]).
  • (K3)
    (Optional) When $\mathcal{Y}$ carries a compatible abelian topological group structure so that translations are defined and continuous, $\mathbf{k}$ is translation-invariant.
Remark.
Assumptions (S1)–(K2) ensure that $\mu_{\mathbf{k}}(P)$ is well-defined for $P\in\mathcal{M}_{b}(\mathcal{Y})$ and that $d_{\mathbf{k}}$ is both discriminative (characteristic) and topologically compatible with weak convergence (metrizing). Assumption (K3) is not needed for the statistical theory, but when satisfied, admits a simple finite-dimensional approximation of the MMD through Fourier-feature representation for fast computation. For example, when $\mathcal{Y}=\mathbb{R}^{d}$, the rate of $d_{\mathbf{k}}$ can be independent of $d$ ([38]); More details are provided in Section A.10 of the Supplementary Material.
Assumptions (S1) and (K1)–(K3) cover many common outcome types, including $\mathcal{Y}=\mathbb{R}^{d}$, graph representations in $0,1]^{m(m-1)/2}$ or $\mathbb{R}^{m\times m}_{\mathrm{sym}}$, torus-valued data $\mathbb{T}^{d}$ (via $\theta\mapsto(\cos\theta,\sin\theta)$), and the simplex $\Delta^{p-1}$, by equipping $\mathcal{Y}$ with standard kernels such as the Gaussian RBF $\mathbf{k}(y,y^{\prime})=\exp\{-\|y-y^{\prime}\|^{2}/(2\sigma^{2})\}$ with $\sigma>0,$ or its restriction after embedding; see Section A.10.
3 SDRF Algorithm and Theoretical Result
Random forests are machine learning methods that combine a CART-style [27] tree ensemble with a bagging step [15], and can also be viewed as kernel-type estimators (see, e.g. [54, 72]). Traditional random forests were designed to estimate the conditional mean regression function and prediction in classification problems [17], and several works have documented their strong empirical performance [26, 31]. More recently, conditional distributional versions of random forests based on kernel mean embeddings have been proposed for i.i.d. data [21]. Our proposed method, SDRF, constructs a kernel estimator $\widehat{\omega}_{\mathbf{x}}(\cdot)$ that directly targets conditional distributions and appropriately incorporates the survey design. Given the estimated neighborhood and weights (which together define $\widehat{\omega}_{\mathbf{x}}(\cdot)$), we can directly construct an estimator of the conditional distribution $\widehat{P}_{\mathcal{S}^{N}\mid\omega}(\mathbf{x})$, and hence a plug-in estimator for $\Psi(\mathbf{x})$, namely $\Psi(\mathbf{x})$: $\Psi(\widehat{P}_{\mathcal{S}^{N}\mid\omega}(\mathbf{x}))$.
We introduce additional notation to help clarify SDRF. Denote $B$ the number of trees, $T_{b}$ the bb–th tree fitted on re-sampled data $D^{*}_{b}=\{(X_{i},Y_{i},1/\pi_{i}\times n^{*}_{b,i})\}_{i=1}^{n_{s}}$, where $n^{*}_{b,i}\geq 0$ are resampling multipliers. Let $q\in(0,1)$ denote the PSU-level selection probability for inclusion in the splitting subsample $D^{*}_{b,\mathrm{split}}$, and let $\mathcal{I}(D^{*}_{b,\mathrm{split}})\subseteq[n_{s}]$ be the set of subject indices contained in $D^{*}_{b,\mathrm{split}}$. Denote $L(\cdot)$ the oracle partition estimated by SDRF that uses the full finite population. This reduces to fitting an RF on i.i.d. data. Denote $L^{*}(\cdot)$ the partition estimated using training data $D^{*}_{train}$ that involves re-sampled weights $n^{*}$. $L(\cdot)$ and $L^{*}(\cdot)$ map any given $\mathbf{x}$ to a region in $\mathcal{X}$. The complete SDRF implementations details are summarized in Algorithm 1. In the following subsections, we discuss in further detail key components, especially those that concern adaptations to design, and introduce their associated theoretical properties.
Algorithm 1 Survey Distributional Random Forest (SDRF)
[…]
0: Random forest ensemble $\mathcal{T}=\{T_{1},\dots,T_{B}\}$
[…]
2:for $b=1,\dots,B$ do
3:  Draw resample $D^{*}_{b}=\{(X_{i},Y_{i},1/\pi_{i}\times n^{*}_{b,i})\}_{i=1}^{n_{s}}$ respecting the design.
4:  Draw $\forall k(i)$, $I(k(i)\in D^{*}_{b,split})\sim Bernoulli(q)$.
5:  Split at the PSU level $D^{*}_{b}=D^{*}_{b,split}\bigcup D^{*}_{b,est}$.
6:  Train decision tree $T_{b}$ on $D^{*}_{b,split}=\{(X_{i},Y_{i},\;1/\pi_{i}\times n^{*}_{b,i}\times 1/q)\}$ by:
7:  Randomly select mtrymtry features.
8:  Choose the split that maximizes $M^{*}(\mathrm{split};D^{*}_{\mathrm{split}})=\frac{N^{*}_{L}N^{*}_{R}}{N^{*2}_{Pa}}\|\mu_{\mathbf{k}}(\hat{P}^{*}_{L})-\mu_{\mathbf{k}}(\hat{P}^{*}_{R})\|^{2}_{\mathcal{H}}$ until reaching maximum depth, minimum node size, or maximum weight ratio.
9:  $\mathcal{T}\leftarrow\mathcal{T}\;\bigcup\;\{T_{b}\}$.
10:end for
11:return For all $b$ with $i\in\mathcal{I}(D^{*}_{b,est})$, $\widehat{\omega}_{i}(\cdot)=\frac{1}{B}\sum_{b=1}^{B}\frac{n^{*}_{bi}/(\pi_{i}(1-q))\,I(X_{i}\in L^{*}_{b}(\mathbf{x}))}{\sum_{i\in\mathcal{I}(D^{*}_{b,est})}n^{*}_{bi}/(\pi_{i}(1-q))\,I(X_{i}\in L^{*}_{b}(\mathbf{x}))}$.
12: For $\Psi(\cdot)$, obtain plug-in estimator $\Psi(\widehat{P}^{N}_{Y\mid X=\mathbf{x}})=\Psi(\sum_{i=1}^{n_{s}}\widehat{\omega}_{i}(\mathbf{x})\;\mathbf{k}(Y_{i},\,\cdot))$.
3.1 Bagging and Honesty
Resampling with Design
Bagging is a key ingredient of random forests [15], but its study is primarily conducted in an i.i.d. setting. In the i.i.d. case, it is well established that bagging introduces diversity across trees and allows the ensemble to approximate complex regression operators [16, 18].
Under a survey design, however, Efron’s classical i.i.d. bootstrap [29] can distort the dependence structure and unequal-probability features induced by the design, and thus may fail to preserve the survey-weighted empirical measure in a controlled way.
We adopt a design–aware bootstrap (BS) method in line with [95] as the default resampling method for SDRF. It generates multipliers $n^{*}$ that determine the additional up– or down– weighting applied to the sampled distribution, and satisfy properties (R1)–(R3) as detailed in Lemma 3.1.
Lemma 3.1.
For single–stage designs including Poisson, SRSWOR, and PPS-with-replacement designs, and two–stage clustering designs with bounded cluster sample-size, possibly with stratification, the pseudo-population bootstrap described in Section A.5.1 of the Supplementary Material in the spirit of [95] (with minor modification) generates unit (or SSU–level) multipliers $\{n_{i}^{*}\}_{i\in[n_{s}]}$ satisfying:
  1. On resample multipliers $n^{*}$
  2. (R1)
    The conditional law of $n^{*}_{i}$ depends on the design only through $\xi^{N}$ and $Z^{N}$ that contains known cluster and strata structure. Each stratum is re–sampled separately, and if the design has two stages, the second stage resampling multipliers are independent across distinct PSUs conditional on first–stage resamples.
  3. (R2)
    $\mathbb{E}(n_{i}^{*}\mid\xi,Z^{N})=1\text{, and }\sup_{i}\mathbb{E}\bigl|n_{i}^{*}\bigr|^{2+\epsilon}<\infty\text{ for some }\epsilon>0.$
  4. (R3)
    $\exists C<\infty$ a constant such that $\sum_{i\neq j}\bigl|Cov(n_{i}^{*},n_{j}^{*}\mid\xi,Z^{N})\bigr|\;\leq\;C\sum_{i}\operatorname{Var}(n_{i}^{*}\mid\xi,Z^{N})\,\,a.s.$ for all $N$.
Remark.
(R1) ensures that the multipliers are generated in a design-adapted and symmetric way and prevents the resampling step from introducing extraneous dependence on unobserved design features. (R2) enforces centering at the obtained sample and stable fluctuations, which are common for BS multipliers. (R3) is a weak dependence requirement across units. It is convenient for deriving concentration bounds while allowing for the existence of dependency.
The method by [95] belongs to the class of pseudo-population-based resampling methods. To summarize: For each re-sample, we first use design information such as $\{w_{i}\}_{i\in[n_{s}]}$ to re-construct via the multinomial distribution a pseudo finite population, with expected size equal to $\widehat{N}=\sum_{i\in[n_{s}]}1/\pi_{i}$; then we re-apply the sampling design $\mathbf{p}$ on the re-constructed finite population to obtain re-samples, thereby respecting the design structure. We refer interested readers to [63, 23] for a review of common resampling methods under design.
Honesty with Design
For random forests algorithms, honesty refers to $D_{est}\mathrel{\perp\!\!\!\perp}D_{split}$, where $D_{split}$ is used for construction of the tree structure $L$ and $D_{est}$ used for populating the leaves with values to form the kernel estimate $\widehat{\omega}(\mathbf{x})$ ([3, 93, 4]). With i.i.d. data, any disjoint partition of the observations yields honesty by construction. However, data from design carries possible dependency across subjects, so separation must be made at the PSU level to ensure the maximum level of independence conditioning on $\xi$. This motivates the introduction of $q\in(0,\,1)$. Alternatively, if one has access to data obtained from a continuous survey under the same design and is willing to assume that $P_{Y|X}$ stays unchanged between two consecutive time points, it is also possible to directly use one sample as $D_{est}$ and the other as $D_{split}$.
In practice, when the number of sampled PSUs is small, the effective sample size (for example, $n_{\text{eff}}=\frac{(\sum_{j}1/\pi_{j})^{2}}{\sum_{i}(1/\pi_{i})^{2}}$ for single–stage designs, [49]) decreases as depth increases. As we discuss later, the performance of the MMD–based split criterion hinges on the effective sample size of $D_{split}$. In this case, one might consider spending the PSUs gradually over the tree growing process. See Algorithm 2, Supplementary Material for details.
3.2 Splitting Rule
The splitting criterion of an RF algorithm guides the formation of tree structure by determining how the feature space should be partitioned into increasingly homogeneous child nodes. Consider any node where a split is possible (has not met the stopping criteria). Denote $P_{a}\subseteq\mathcal{X}$ the feature space of that node. A candidate split can be parametrized as $\theta=(j,\;t)\in\Pi$, $j=1,\dots,p$, $t\in Range(X_{j})\subseteq[a_{j},\;b_{j}]$. This partitions $P_{a}$ into child regions $C_{L}(\theta):=\{x\in P_{a}:x_{j}\leq t\}$ and $C_{R}(\theta):=\{x\in P_{a}:\;x_{j}>t\}$. With RF considered as an adaptive kernel estimation algorithm, given any fixed region $L$, we can estimate the conditional probability via Hájek–type estimator. We find the split by maximizing the MMD between the KMEs of the distribution estimates between the two child regions. Formally, $\hat{\theta}=\arg\max_{\theta\in\Pi}M^{*}_{n_{s}}(\theta)$, where
$\begin{split}M^{*}_{n_{s}}(\theta)=\frac{\hat{N}_{L}\hat{N}_{R}}{\hat{N}^{2}_{P_{a}}}\,\left\|\mu_{\mathbf{k}}\left(\frac{1}{\hat{N}_{L}}\sum_{\{i\in[n_{s}]:X_{i}\in C_{L}\}}\frac{n^{*}_{b,i}}{q\times\pi_{i}}\;\delta_{Y_{i}}\right)-\mu_{\mathbf{k}}\left(\frac{1}{\hat{N}_{R}}\sum_{\{i\in[n_{s}]:X_{i}\in C_{R}\}}\frac{n^{*}_{b,i}}{q\times\pi_{i}}\;\delta_{Y_{i}}\right)\right\|^{2}_{\mathcal{H}},\end{split}$ (1)
with $\hat{N}_{\ell}=\sum_{i\in[n_{s}]}I(X_{i}\in C_{\ell})\frac{n^{*}_{b,i}}{q\times\pi_{i}}$, $\ell\in\{L,R\}$. When under stratified two–stage designs, notice that the same form for $M^{*}_{n_{s}}$ holds, except that we replace the sum over subjects by sum $\sum_{h,k}\sum_{i\in\mathcal{U}_{hk}}$ over SSUs $(h,k)$ and then over subjects within SSUs, if exists. We first provide a theorem justifying the choice of this split criterion.
Theorem 3.2 (Local design/model consistency of the MMD split).
Fix a parent region $P_{a}\subset\mathcal{X}$ at depth $d$ and consider axis-aligned potential splits $\theta=(j,t)\in\Pi:=\{(j,t):j\in[p],\ t\in[a_{j},b_{j}]\}$ which defines child regions $C_{L}(\theta)$ and $C_{R}(\theta)$. Let $M^{*}_{n_{s}}(\theta)$ denote the (design-resampled) split score in (1), and let $M^{N}(\theta)$ be its finite-population analog. Suppose (D1)–(D4), (R1)–(R3), (S1), (K1)–(K2), and (A2)–(A3) described in Section 3.3. Suppose further $\mathcal{X}\subset\mathbb{R}^{p}$ is bounded. Then, conditional on the realized finite population,
$\sup_{\theta\in\Pi}\bigl|M^{*}_{n_{s}}(\theta)-M^{N}(\theta)\bigr|\xrightarrow{p_{d}}0.$
If, in addition, the finite population is an i.i.d. realization from a super-population law $P$ and $M(\theta)$ denotes the corresponding population score, then $\sup_{\theta\in\Pi}\bigl|M^{*}_{n_{s}}(\theta)-M(\theta)\bigr|\xrightarrow{p_{\otimes}}0.$ Moreover, if $M^{N}(\theta)$ has a unique maximizer $\theta_{0}$ and the splits are chosen over a random set of candidates $\Pi(U)$ induced by mtrymtry features such that $\mathbb{E}_{U}[\mathrm{dist}(\theta_{0},\Pi(U))]\to 0$, then $M^{N}$ satisfies the modulus of continuity condition with $d(\theta,\theta^{\prime}):=d((j,t),(j^{\prime},t^{\prime}))=1(j\neq j^{\prime})+|t-t^{\prime}|$, and any maximizer $\hat{\theta}_{n_{s}}\in\arg\max_{\theta\in\Pi(U)}M^{*}_{n_{s}}(\theta)$ satisfies $\hat{\theta}_{n_{s}}\to_{p_{d}}\theta_{0}$ (and hence also $\hat{\theta}_{n_{s}}\to_{p_{\otimes}}\theta_{0}$ under the super-population model).
Remark.
Theorem 3.2 provides a local result: for a fixed parent node $P_{a}$, the empirical MMD criterion $M^{*}_{n_{s}}(\theta)$ uniformly approximates its finite-population oracle counterpart $M^{N}(\theta)$ and also the super-population score $M(\theta)$. The uniformity over $\theta\in\Pi$ is essential for split selection, as it permits an argmax continuous-mapping argument and yields consistency of the selected split $\hat{\theta}_{n_{s}}$ whenever the oracle maximizer $\theta_{0}$ is unique ([90]). The condition $\mathbb{E}_{U}[\mathrm{dist}(\theta_{0},\Pi(U))]\to 0$ makes explicit that random feature subsampling (via mtrymtry) does not distort the split choice asymptotically, provided the candidate set becomes sufficiently dense around $\theta_{0}$ in expectation. Although extending from a fixed node to an entire data-adaptive tree requires additional control of the growing process, this theorem sheds light on why one might believe the partitions found via resampled data to be close to that if using the full finite population, which is in particular assumption (B2) of Theorem 3.4 required for consistency in MMD metric.
One might wonder that, since resamples are determined by non–negative multipliers $n^{*}$, if one can use an averaged multiplier from $M$ (conditional) independent draws to form one re-sample. This can help stabilize the resample, especially when the original sampling weights vary greatly, but can also reduce diversity across trees. Proposition 3.3 below provides a detailed characterization of its consequence.
Proposition 3.3 (Effect of averaging $M$ resampling draws on split-score stability).
Fix a node $P_{a}$ and a candidate split $\theta$ that induces children $C_{L},C_{R}$ with $N_{P_{a}}:=|\{i\leq N:X_{i}\in P_{a}\}|$, $N_{\ell}:=|\{i\leq N:X_{i}\in C_{\ell}\}|$. Suppose there exists $\alpha\in(0,1/2]$ such that $\frac{N_{L}}{N_{P_{a}}},\ \frac{N_{R}}{N_{P_{a}}}\in[\alpha,1-\alpha].$ Let $M_{N}(\theta)$ be the finite-population split score, $M_{n_{s}}(\theta)$ its design-weighted sample analogue, and $M_{n_{s}}^{*(M)}(\theta)$ the resampled score computed using averaged multipliers $\bar{n}_{i}^{*}:=M^{-1}\sum_{b=1}^{M}n_{b,i}^{*}$ (equivalently $\bar{\delta}_{i}:=\bar{n}_{i}^{*}-1$). Under (D1)–(D4), (R1)–(R3), and (K1)–(K2), if the multiplier increments $\delta_{b,i}:=n_{b,i}^{*}-1$ are conditionally sub-gamma and negatively dependent given $\xi$, which are satisfied by the methods described in Section A.5.1, then
$\bigl|M_{n_{s}}^{*(M)}(\theta)-M_{N}(\theta)\bigr|=O_{p}\!\left(N_{P_{a}}^{-1/2}\right)\;+\;O_{p}\!\left(\frac{1}{\sqrt{M}}\sum_{\ell\in\{L,R\}}n_{\mathrm{eff},\ell}^{-3/2}\right),$
where $n_{\mathrm{eff},\ell}$ is the design-weighted effective sample size ([49]) in child $\ell$ induced by $\theta$.
Remark.
Proposition 3.3 separates two sources of split-score fluctuation: an intrinsic design term $|M_{n_{s}}(\theta)-M_{N}(\theta)|=O_{p}(N_{P_{a}}^{-1/2})$ which concerns the sampling noise already present even with fixed $n^{*}$, and a resampling term $|M_{n_{s}}^{*(M)}(\theta)-M_{n_{s}}(\theta)|=O_{p}\!\big(M^{-1/2}\sum_{\ell\in\{L,R\}}n_{\mathrm{eff},\ell}^{-3/2}\big)$.
The cross-over threshold occurs when these terms balance: if $M\gtrsim\tilde{M}$ where $\tilde{M}\asymp N_{P_{a}}\!\big(\sum_{\ell\in\{L,R\}}\\ n_{\mathrm{eff},\ell}^{-3/2}\big)^{2},$ then the design term dominates and increasing $M$ further yields negligible stability gains. At deep nodes (small $n_{\mathrm{eff},\ell}$), this threshold is reached quickly: for instance, if both children have $n_{\mathrm{eff},\ell}\approx 100$ and $N_{P_{a}}\approx 200$, then $\tilde{M}\approx 8$ suffices. Beyond this point, split stability is governed by the effective sample sizes—controlled via minimum-node-size constraints, shallower depths, or larger subsampling fractions qq—rather than by the number of resamples.
Moreover, the marginal gain from increasing $M$ deteriorates rapidly: doubling from $M$ to $2M$ reduces the resampling component by only $2^{-1/2}\approx 0.71$, and more generally $(1+\Delta/M)^{-1/2}\approx 1-\Delta/(2M)$ when $\Delta\ll M$. Thus, in typical regimes where $M$ moderately exceeds $\tilde{M}$, averaging offers limited benefit and computational resources are better invested in increasing $n_{\mathrm{eff},\ell}$.
3.3 Prediction and Aggregation
We now establish the main result of the paper, namely the consistency of the proposed SDRF plug-in estimator $\widehat{\Psi}_{N}(\mathbf{x})$. The argument proceeds in two steps. First, Theorem 3.4 shows that, under the additional algorithmic conditions (A1)–(A3), (B1)–(B2) detailed below, the MMD between the SDRF estimate of the conditional distribution and the corresponding target conditional distribution converges to zero in $P_{\mathcal{S}^{N}\mid\omega}$-probability. Second, Corollary 3.5 shows that this distributional convergence implies the desired consistency of $\widehat{\Psi}_{N}(\mathbf{x})$ for $\Psi(\mathbf{x})$ in both the model- and design-based senses. The implication follows because SDRF is constructed to directly separate conditional distributions, and because $\Psi$ is continuous as a mapping on $\mathcal{M}_{b}(\mathcal{Y})$.
Theorem 3.4.
Suppose (S1) for the space $\mathcal{Y}$, (D1)–(D4) for the survey design $\mathbf{p}$, (R1)–(R3) for the resample multiplier $n^{*}$, and (K1)–(K3) for the reproducing kernel $\mathbf{k}$. Suppose for the true KME of the conditional measure in the model–space, $\exists L_{lip}>0$ such that for any $\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{X}\subseteq\mathbb{R}^{p}$, $\left\|\mathbb{E}_{P}[\mathbf{k}(Y,\cdot)\mid X=\mathbf{x}]-\mathbb{E}_{P}[\mathbf{k}(Y,\cdot)\mid X=\mathbf{x}^{\prime}]\right\|_{\mathcal{H}}\leq L_{lip}\|\mathbf{x}-\mathbf{x}^{\prime}\|$. Suppose also the following:
  • Algorithm regularization
  • (A1)
    Shape regularity: For a.s. $L(\mathbf{x})$ for any $\mathbf{x}$ in the sampled covariate space satisfying $diam(L(\mathbf{x}))<k_{N}$,  $\exists\;0<c_{1}\leq 1$ s.t. $\bar{B}_{r_{N}}(\mathbf{x})\subseteq L(\mathbf{x})$, $r_{N}=c_{1}k_{N}$.
  • (A2)
    There exist constants $0<m<M<\infty$ such that for all radius r, of some $L(\mathbf{x})$ for the given $\mathbf{x}$ such that $0<r\leq k_{N}$, $mr^{p}\leq\#\{i:X_{i}\in B_{r}(\mathbf{x})\}/N\leq Mr^{p}$, where $B_{r}(\mathbf{x})$ the ball centered at $\mathbf{x}$ with radius $r$.
  • (A3)
    For any $L^{*}$ found by $D_{split}$, when populating the values with $D_{est}$, all leaves receive at least $n_{L^{*},min}$ subjects. In other words, $n_{L^{*}(\mathbf{x})}=\sum_{i}^{N}\xi_{i}I(X_{i}\in L^{*}(\mathbf{x}))\geq n_{L^{*},min}>0$.
  • Algorithmic behavior
  • (B1)
    For any $\mathbf{x}$ and some deterministic chosen sequence $k_{N}\to 0$ yet $k^{p}_{N}N\to\infty$ as $N\to\infty$, there exists $t_{N}\to 0$ such that $P_{L}(diam(L(\mathbf{x}))\geq k_{N})\leq t_{N}$, where $iam(L(\mathbf{x}))=\sup_{x_{1},x_{2}\in L(\mathbf{x})}d(x_{1},x_{2}), dd$ some metric on $\mathcal{X}^{p}$.
  • (B2)
    $\sup_{x\in\mathcal{X}_{1}\cup\mathcal{X}_{2}}\mathbb{P}\{X\in L^{*}(x)\,\Delta\,L(x)\}=o_{p}(d_{N})$ for some $d_{N}=o(k_{N}^{p})$, where $\Delta$ denotes symmetric difference: $A\Delta B:=(A\setminus B)\;\cup\;(B\setminus A)$.
Denote $\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x})$ the estimate from SDRF. Then $d_{\mathbf{k}}(\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}),{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}))\to_{p_{d}}0$ as $N,B\to\infty$.
Remark.
Regularization–type conditions (A1)–(A3) can be controlled via adaptive stopping criteria. (A1) imposes a geometric control on the shape of $L$ so that they do not be stretched. (A2) requires a local mass control for the local empirical density to be well-behaved. (A3) ensures the local density estimate stays bounded. (B1)-(B3) formalize how the partitions produced by the algorithm behave as the finite population size grows. (B1) is a shrinking diameter assumption that mirrors the adaptive concentration result proved for i.i.d. forests ([94]). (B2) requires that the thickness of the error band be negligible relative to the leaf’s own diameter. In other words, the rate at which the leaf $L^{*}(\cdot)$ function estimated from the bootstrap sample approaches the leaf $L(\cdot)$ estimated if using the full finite population must be faster than the oracle leaf’s own shrinking-volume rate. Under the required relationship between $d_{N}$ and $k_{N}$, together with (A1) regularity conditions on the geometry, the symmetric-difference region $L^{*}(x)\triangle L(x)$ has measure $O_{p}(d_{N}/k_{N}^{p})$, which implies that both the tree estimator $\widehat{T}^{*}_{N}(L_{\xi}^{*}(x))$ and its conditional variance coincide with their oracle counterparts up to $o_{p}(N^{-1/2})$ and $o_{p}(N^{-1})$, respectively. Among all the assumptions, (B2) is the hardest to directly control. Thus we provided Theorem 3.2 that partially shed light on conditions under which (B2) can be reasonably assumed to hold.
Corollary 3.5.
Suppose $\{(X_{i},Y_{i},Z_{i})\}_{i=1}^{N}$ i.i.d. realizations from $P_{\Omega}(\cdot)$. Suppose $\mathbf{k}$ satisfying (K1)–(K3), and $\mathcal{Y}$ locally compact Polish (satisfying (S1)). Denote ${P}_{\mathcal{S}^{N}|\omega}(\mathbf{x})$ the true conditional measure of $Y|X=\mathbf{x}$ in the design space and $\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x})$ an estimate for it. Denote $P_{\omega}(\mathbf{x})$ the true conditional measure of $Y|X=\mathbf{x}$ in the super population space. Then for any specified $\mathbf{x}\in\mathcal{X}$ and $\Psi:(\mathcal{M}_{b}(\mathcal{Y}),\;d_{\mathbf{k}})\to(\mathcal{V},d)$ a measurable functional continuous at $P^{N}(Y|X=\mathbf{x})$ for a.e. $\omega\in\Omega$, if   $d_{\mathbf{k}}(\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}),{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}))\to_{p_{d}}0,\text{ then }$
$\Psi(\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}))$ $\to_{p_{d}}\Psi({P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}))\;\;\text{ (design consistency) },\text{ and }$
[…]
Remark.
Corollary 3.5 reduces design– and model–based consistency of $\widehat{\Psi}_{N}(\mathbf{x})$ to control of $\widehat{P}_{S^{N}\mid\omega}(\mathbf{x})$ in the $d_{\mathbf{k}}$-metric, together with continuity of the functional $\Psi$ at the relevant target law. Many standard summaries satisfy this requirement under mild regularity. For example, when $\mathcal{Y}=\mathbb{R}^{d}$, linear functionals such as conditional means $\Psi(P)=\int y\,dP(y)$ and, more generally, $\Psi_{g}(P)=\int g(y)\,dP(y)$ for $g\in C_{b}(\mathcal{Y})$ are continuous under weak convergence. For distributional functionals, $\Psi_{t}(P)=F_{P}(t)$ and $\Psi_{t}(P)=P(Y>t)$ are continuous at $P$ whenever $P(\{t\})=0$. The quantile functional $\Psi_{q}(P)=F_{P}^{-1}(q)$ is continuous at $P$ whenever the $q$-quantile is unique. Equivalently, that $F_{P}$ has no flat segment at level $q$. The kernel mean element itself and its norm $\Psi(P)=\|\mu_{\mathbf{k}}(P)\|_{\mathcal{H}}$, and the MMD between the true conditional $P$ to a reference distribution $Q$, $d_{\mathbf{k}}(P,Q)$, are also examples.
3.4 Implementation Considerations
We collect practical implementation considerations for SDRF, consolidating several points noted earlier.
  • For the re–sampling, ([22]) showed that asymptotically, generating one pseudo population and then drawing $B$ samples from it (equivalently, realizing $B$ times $n^{*}$ conditioning on the realized pseudo population) is asymptotically equivalent to realizing independently $B$ pseudo population “and” the re-sample.
  • RBF–type kernel is a common class of kernels that is translation–invariant. Median-heuristic, i.e., specifying the bandwidth $\sigma>0$ to be the median of pairwise sampled distance, is a commonly adopted, and reasonable choice ([35, 64]).
  • Hyperparameters governing the stopping criteria are introduced to help ensure regularity-type conditions (A1)–(A3). For finite samples, it may be preferable to use a moderate number of shallow trees to keep local sample weights well-controlled. Moreover, when the number of PSUs is large but PSU sizes are small, one may consider omitting second-stage resampling (Section A.5.1).
  • Because SDRF yields a kernel-type estimator, one could further post-process $\widehat{P}(Y\mid X=\mathbf{x})$ (or $\widehat{\Psi}(\mathbf{x})$) with an additional kernel-smoothing step [92] in $\mathbf{x}$ to reduce finite-sample variability and stabilize the resulting estimates, particularly when it is reasonable to assume that the underlying target varies smoothly in $\mathbf{x}$.
To conclude, up to this point, we have illustrated the necessity of an appropriate method when data arise under sampling designs as opposed to being i.i.d.; described in detail our proposed algorithm as a design-aware solution; proved both types of consistency for the resulting estimator under reasonable assumptions; discussed practical implementation-side considerations; and outlined specific example targets for which this algorithm applies. The algorithm is flexible in the choice of $\Psi$, the space of $Y$, as well as the type of design.
4 Simulations
4.1 Data Generation
To assess the finite-sample performance of the proposed survey-calibrated distributional random forest (SDRF), we conduct a simulation study designed to mimic the NHANES case study that follows a complex multistage survey design. Data are generated under a stratified cluster sampling scheme with probability-proportional-to-size (PPS) selection of primary sampling units (PSUs), followed by simple random sampling without replacement (SRSWOR) of secondary units within selected PSUs.
Our goals are twofold:
  1. 1.
    Quantify gains in statistical accuracy and efficiency as the sample size increases and compare SDRF with DRF [21], an existing distributional random forest that does not incorporate survey weights in the estimation process.
  2. 2.
    Study the effect of the number of trees, the sample size and the number of PSUs per stratum $N_{\mathrm{psu},h}$, on the finite-sample performance.
We begin by defining the generative model in the super–population framework. For each unit $i=1,\dots,N$, let
$Y_{i}=(Y_{i1},Y_{i2})^{\top}\in\mathbb{R}^{2},\qquad X_{i}=(X_{i1},\dots,X_{ip})^{\top}\in\mathbb{R}^{p},\qquad Z_{i}\in\mathbb{R},$
where $Y_{i}$ is the outcome vector, $X_{i}$ the predictor vector and $Z_{i}$ a design variable that influences sampling but is not observed at the unit level in the analysis.
The finite population $\{(Y_{i},X_{i},Z_{i}):i=1,\dots,N\}$ is generated as follows: For each $i$,
$Z_{i}\overset{\mathrm{iid}}{\sim}\mathcal{N}(0,1),\quad X_{i1}\overset{\mathrm{iid}}{\sim}\text{Uniform}(0,1),$
$B_{i}\overset{\mathrm{iid}}{\sim}\operatorname{Bernoulli}\!\left(\tfrac{1}{2}\right),\quad X_{i2}=2B_{i}-1,$
$X_{ij}\overset{\mathrm{iid}}{\sim}\mathcal{N}(0,1),\quad j=3,\dots,p,$
where all variables in the construction above are mutually independent across both units $i$ and indices $j$, unless otherwise stated. We then define the conditional mean of $Y_{i}$ given $(X_{i},Z_{i})$ by
$Y_{i}\;=\;\mu(X_{i},Z_{i})+\Sigma^{1/2}\,\varepsilon_{i},\qquad\varepsilon_{i}\overset{\mathrm{iid}}{\sim}\mathcal{N}(\mathbf{0},I_{2}),\text{ where }$
$\mu(X_{i},Z_{i})=\begin{pmatrix}\mu_{1}(X_{i},Z_{i})\\[2.0pt] \mu_{2}(X_{i},Z_{i})\end{pmatrix},\quad\Sigma=\begin{pmatrix}1&0.3\\[2.0pt] 0.3&1\end{pmatrix},\text{ and }$
$\mu_{1}(X_{i},Z_{i})=1.5\,X_{i2}+(2+50Z_{i})\,X_{i1},\qquad\mu_{2}(X_{i},Z_{i})=1.5\,X_{i2}-Z_{i}\,(3X_{i2}+1).$
$\{\varepsilon_{i}\}^{N}_{i=1}$ are generated independently of $\{(X_{i},Z_{i})\}$. Equivalently, $Y_{i}\mid(X_{i},Z_{i})\sim\mathcal{N}_{2}\bigl(\mu(X_{i},Z_{i}),\,\Sigma\bigr)$ i.i.d. across i.
We now describe the sampling design. Let $H\geq 2$ be the number of strata, and introduce a stratum indicator
$\mathrm{Strat}_{i}\in\{1,\dots,H\},\qquad i=1,\dots,N.$
Each unit is assigned independently to a stratum with equal probability:
$\bigl(\mathds{1}\{\mathrm{Strat}_{i}=1\},\ldots,\mathds{1}\{\mathrm{Strat}_{i}=H\}\bigr)\sim\operatorname{Multinomial}\!\left(1;\tfrac{1}{H},\ldots,\tfrac{1}{H}\right).$
For each stratum $h\in\{1,\dots,H\}$, let
$U_{h}=\{i:\mathrm{Strat}_{i}=h\},\qquad N_{h}=|U_{h}|$
denote the set and size of units in stratum $h$. Within stratum $h$, define the membership map $g_{h}:U_{h}\to\{1,\ldots,N_{h}^{\mathrm{PSU}}\}$ by ordering units in $U_{h}$ by $X_{i1}$ (and breaking ties arbitrarily) and grouping the ordered list into $N_{h}^{\mathrm{PSU}}$ consecutive bins of approximately equal size. Write $g_{h}(i)=j$ if unit $i$ is assigned to PSU $j$, and define
$\mathrm{PSU}_{hj}:=\{i\in U_{h}:g_{h}(i)=j\},\qquad j=1,\ldots,N_{h}^{\mathrm{PSU}}.$
We define measure-of-size as a function of $Z$: For each PSU, $\mathrm{PSU}_{hj}$,
$M_{hj}\;=\;8\mathds{1}(Z_{hj}>0)+2\mathds{1}(\bar{Z}_{hj}\leq 0),\quad\text{where }\bar{Z}_{hj}\;=\;\frac{1}{N_{hj}}\sum_{i\in\mathrm{PSU}_{hj}}Z_{i}$
the PSU-level mean of $Z_{i}$, $N_{hj}=|\mathrm{PSU}_{hj}|$ the number of subjects within that PSU.
First-stage sampling selects PSUs within each stratum $h$ with probabilities proportional to size. Fix an expected number $n_{h}^{(1)}$ of PSUs to be selected in stratum $h$, and define the inclusion probabilities of the first-stage
$\pi^{(1)}_{hj}\;=\;\min\!\Bigl(1,\;n^{(1)}_{h}\,\frac{M_{hj}}{\sum_{j^{\prime}=1}^{N_{h}^{\mathrm{PSU}}}M_{hj^{\prime}}}\Bigr),\qquad j=1,\dots,N_{h}^{\mathrm{PSU}}.$
We then select PSUs by unequal-probability systematic sampling,
$\{\mathrm{PSU}_{hj}:\xi^{(1)}_{hj}=1\}\sim\operatorname{UP\text{-}systematic}\!\bigl(\{\pi^{(1)}_{hj}\}_{j=1}^{N_{h}^{\mathrm{PSU}}}\bigr),$
where $\xi^{(1)}_{hj}\in\{0,1\}$ is the indicator that PSU $j$ in stratum $h$ is selected at the first stage.
In the second stage, conditional on $\xi^{(1)}_{hj}=1$, we sample units within PSU $\mathrm{PSU}_{hj}$ by SRSWOR. Specifically, we select
$n^{(2)}_{hj}\;=\;\lceil 0.3\,N_{hj}\rceil$
units without replacement in each selected PSU. Thus, for any unit $i\in\mathrm{PSU}_{hj}$,
$\pi^{(2)}_{hji}\;=\;\Pr(\xi^{(2)}_{hji}=1\mid\xi^{(1)}_{hj}=1)\;=\;\frac{n^{(2)}_{hj}}{N_{hj}}\;=\;0.3,$
where $\xi^{(2)}_{hji}\in\{0,1\}$ is the second-stage selection indicator.
Under this two-stage design, the overall unit inclusion indicator is $\xi_{hji}=\xi^{(1)}_{hj}\,\xi^{(2)}_{hji}$, with unit inclusion probability $\pi_{hji}=\Pr(\xi_{hji}=1)=\pi^{(1)}_{hj}\,\pi^{(2)}_{hji}$, and corresponding survey weights $w_{hji}=\frac{1}{\pi_{hji}}.$ This completes the generation of analysis dataset $D=\{(X_{i},Y_{i},w_{hji})\}_{\{i:\xi_{hji}=1\}}$. The sampling scheme considered here satisfies the conditional non-informativeness requirement (D1).
4.2 Computation Details
We run the SDRF and DRF ([21]) algorithms with the above data generating process for each of the finite-population size $N\in\{4000,5000,15000,22500\}$ and the number of trees $B\in\{10,30,70,200\}$. For each of the scenario, 200 seeds were run. We let $p=3$ to align with the case study (Section 5). As $N$ increases, we also increase the total number of strata $H$ and the number of PSUs per stratum $N_{\mathrm{psu},h}$. The expected sample size per stratum (and hence the overall sample size) increases with $N$. The sampling mechanism (PPS at stage 1 and SRSWOR at stage 2 with fixed rates) is kept fixed in the values of $N$, which is consistent with the asymptotics of the increase in the population survey.
For analytical evaluation, we report (a) the maximum mean discrepancy (MMD) between the estimated kernel mean embedding and the population kernel mean embedding for the conditional distribution on a hold-out set, and (b) the accuracy in the estimation for the functional
$\Psi(\mathbf{x}=(x_{1},\dots,x_{p}))\;=\;\mathbb{E}\!\left[\,Y_{1}\mid X=\mathbf{x}_{1}\,\right],$
in terms of mean squared error (MSE) of $\widehat{\Psi}(\mathbf{x}_{1})$.
Unlike the training data, which are the complex-sample draws produced by the design, the testing data are generated directly from the super-population model; thus no design information (weights or $Z$) is required to evaluate $\Psi(\mathbf{x}_{1})$ in the test set. We focus on $\Psi(\mathbf{x}_{1})$ to enable visual inspection of pointwise performance across the range of $X_{1}$.
For both random-forest estimators, we use a Gaussian kernel on the outcome space with bandwidth set by the median of pairwise Euclidean distances between the observed $Y$’s. The maximum tree depth is fixed at $8$. The minimum node size is set to $\max\!\left\{\,20,\;\left\lceil\sqrt{n_{s}}\,\right\rceil\right\}$, where $n$ denotes the size of the training-sample. We cap effective sampling weights by $\lambda_{\max}\;=\;5.5\times\frac{\max_{i}w_{i}}{\min_{i}w_{i}}$, where $w_{i}$ are the survey weights. Although additional hyperparameter tuning could further improve performance, we keep the settings minimal because the untuned configurations already reveal the main trends across $N$ and $B$.
We do not compare against alternative survey-calibrated distributional methods beyond SDRF because, to our knowledge, such methods are not currently available in the survey-sampling literature. (The unweighted DRF serves as the i.i.d. baseline and is a logical competitor). All code is implemented under R, version 4.4.2 [73].
4.3 Results
Table 1 reports the mean (SD) of the MMD between the true finite population conditional distribution $P(Y\mid X)$ and its estimate $\widehat{P}(Y\mid X)$ from SDRF. The mean MMD decreases as the finite population size $N$ increases, while increasing the number of trees $B$ primarily reduces the SD (stabilizing forest performance) with comparatively smaller improvements in the mean and diminishing returns beyond moderate $B$. Overall, these finite-sample patterns are consistent with the convergence behavior asserted in Theorem 3.4.
Table 1: Mean (SD) of $100\times d_{\mathbf{k}}(\widehat{P}(Y\mid X),\;P^{N}(Y\mid X))$, where $\widehat{P}(Y\mid X)$ is the distribution estimate from SDRF, across finite population size $N$ and number of trees $B$.
NN $B=10$ $B=30$ $B=70$ $B=200$
4,000 19.60 (2.52) 18.60 (2.27) 18.30 (2.19) 18.30 (2.10)
5,000 19.30 (2.49) 18.40 (2.14) 18.00 (2.14) 18.10 (2.07)
15,000 17.30 (1.86) 16.80 (1.44) 16.70 (1.30) 16.80 (1.23)
22,500 17.40 (1.52) 16.80 (1.23) 16.60 (1.16) 16.70 (1.08)
Table 2: Root mean squared error (RMSE) of the conditional mean estimator $\widehat{\mathbb{E}}[Y_{1}|X_{1}]$ averaged over $X_{1}$ for Survey-DRF (SDRF) and benchmark DRF algorithms as $N$ and the number of trees grow.
NN $N_{\text{PSU},h}$ $\bar{n}_{s}$ Algorithm Number of Trees $B$
10 30 70 200
4,000 80 320 DRF 7.80 7.49 7.32 7.33
SDRF 4.11 3.44 3.29 3.27
5,000 100 400 DRF 8.04 7.55 7.40 7.33
SDRF 3.69 3.16 2.95 2.96
15,000 200 1,200 DRF 7.64 7.45 7.45 7.44
SDRF 2.48 2.11 2.00 1.95
22,500 300 1,800 DRF 7.52 7.37 7.33 7.32
SDRF 2.16 1.83 1.74 1.74
Figure 1: Pointwise mean square error (MSE) of the estimator $\widehat{\mathbb{E}}[Y_{1}|X_{1}=\mathbf{x}_{1}]$ from SDRF and DRF. Each point represents an average over 200 seeds.
Figure 2: Pointwise standard error (SD) of the estimator $\widehat{\mathbb{E}}[Y_{1}|X_{1}=\mathbf{x}]$ from SDRF and DRF. Each point represents an average over 200 seeds.
Table 2 and Figures 12 compare SDRF with the i.i.d.-based DRF for estimating the regression functional $\Psi(\mathbf{x}_{1})=\mathbb{E}[Y_{1}\mid X_{1}=\mathbf{x}_{1}]$. Averaged over a uniform grid of 200 points in $[0,1]$, SDRF achieves smaller RMSE than DRF for all finite population sizes $N$ and ensemble sizes $B$ considered (Table 2).
Figure 1 shows the corresponding pointwise MSE: for both methods, MSE decreases with $B$, but DRF exhibits a clear error floor that persists as $N$ and $B$ increase, whereas SDRF’s MSE decreases uniformly towards the truth, including in the upper range of $X_{1}$ where sampling weights are more variable. Figure 2, combined with Figure 1, suggests that the MSE gap is driven primarily by persistent bias in DRF. This is expected as reasoned in Section 1. Also, Figure 2 shows that pointwise variability decreases with $B$ for both methods and remains comparable between SDRF and DRF; notably, SDRF does not exhibit an appreciable variance inflation despite the additional randomness introduced by its design-aware resampling step.
5 Biomedical Public Health Case Study
This section demonstrates the practical relevance of survey-calibrated distributional random forests (SDRF) for performing data analysis tasks in nationally representative studies. We use data from the National Health and Nutrition Examination Survey (NHANES), which employs a multistage, stratified, cluster sampling design to represent the U.S. civilian, noninstitutionalized population: primary sampling units (PSUs) are selected at the county level, followed by smaller geographic areas and then individuals. All analyzes performed incorporate survey weights and design features provided by the Centers for Disease Control and Prevention (CDC) [^1] https://wwwn.cdc.gov/nchs/nhanes/Default.aspx . More specifically, in this paper, we focus on the 2011–2012 NHANES cycle and apply an SDRF algorithm that sampling mechanism of the NHANES study during bagging/subsampling and throughout tree construction (e.g. split selection and node estimation).
The objective of the statistical analysis is to characterize the conditional joint distribution of two glucose biomarkers used to diagnose and monitor the progression of diabetes, fasting plasma glucose (FPG) and glycated hemoglobin (HbA1c), given demographic and anthropometric covariates: sex, age, and body mass index (BMI). Our primary clinical goal is to quantify how the population distribution of this bivariate outcome varies between these characteristics, to support the detection of diabetes, the evaluation of glycemic-control, and healthcare planning through the trends observed in the subgroups of the population. A secondary clinical objective is to evaluate whether the observed heterogeneity motivates individual diabetes diagnostic thresholds for FPG and HbA1c, in contrast to fixed cutoffs proposed by medical associations such as the American Diabetes Association (ADA) [2], which ignore specific individual characteristics. For analytical purposes, a total of $N=2732$ individuals with and without diabetes were included in the analysis.
Let us go to define mathematically the problem. Let the clinical outcome be $Y=(Y_{1},Y_{2})^{\top}$, where $Y_{1}$ denotes fasting plasma glucose (FPG) and $Y_{2}$ denotes HbA1c. Let the covariate vector be $X=(X_{1},X_{2},X_{3})^{\top}$, where $X_{1}=\mathrm{Age}$ and $X_{2}=\mathrm{BMI}$ are continuous, and $X_{3}=\mathrm{Sex}\in\{0,1\}$ is binary with $0=\text{male}$ and $1=\text{female}$.
From an empirical perspective, we run SDRF and estimate the conditional distribution of the bivariate result $Y=(\mathrm{FPG},\mathrm{HbA1c})^{\top}$ given covariates $X=\mathbf{x}$. Specifically, we computed the estimated conditional CDF
$\hat{F}(y\mid\mathbf{x})=\hat{P}^{N}_{\mathrm{HT}}\big(Y\leq y\mid X=\mathbf{x}\big),\qquad\text{for all }y\in\mathcal{Y},\ \mathbf{x}\in\mathcal{X},$
where inequality $Y\leq y$ is understood componentwise. For representation purposes, we evaluate $\hat{F}(\cdot\mid\mathbf{x})$ on a grid of covariate combinations:
$\mathrm{Age}\in\{30,50,70\},\quad\mathrm{BMI}\in\{20,30,40\},\quad\mathrm{Sex}\in\{0,1\}.$
Let $\hat{\omega}_{i}(x)$ denote the SDRF distribution-regression weights for observation $i\in[N]$ at covariate value $x\in\mathcal{X}$. Using these weights, we compute the estimated conditional mean and covariance,
$\hat{\mu}(x)=\sum_{i=1}^{N}\hat{\omega}_{i}(x)\,Y_{i}\quad\text{and}\quad\hat{\Sigma}(x)=\sum_{i=1}^{N}\hat{\omega}_{i}(x)\,\bigl(Y_{i}-\hat{\mu}(x)\bigr)\bigl(Y_{i}-\hat{\mu}(x)\bigr)^{\top}.$
We then define a conditional Mahalanobis score (weighted) for any pair $(x,y)\in\mathcal{X}\times\mathcal{Y}$ by
$S(x,y)=\bigl(y-\hat{\mu}(x)\bigr)^{\top}\hat{\Sigma}(x)^{-1}\bigl(y-\hat{\mu}(x)\bigr).$
Let $\{(x_{a},y_{a})\}_{a=1}^{N}$ denote the observed random sample and let $\hat{q}_{1-\alpha}$ be the empirical $(1-\alpha)$ quantile of the sample $\{S(x_{a},y_{a})\}_{a=1}^{N}$ (using the survey weights).
In the spirit of [53] (for finite-dimensional Euclidean responses) and [57] (for metric space outcomes), we define the tolerance region $(1-\alpha)$ at the covariate value $x\in\mathcal{X}$ as
$\hat{C}^{\alpha}(x)=\bigl\{\,y\in\mathcal{Y}:\;S(x,y)\leq\hat{q}_{1-\alpha}\,\bigr\}.$
Compared to the natural competitor [53], our approach is more flexible: it requires a single model, avoids user-chosen smoothing parameters due to the use of multiple local polynomial regression, and enables valid inference under the complex design of the NHANES survey.
Figures 3 and 4 show tolerance regions (in expectation) for $\alpha\in\{0.10,0.25,0.50,0.75,0.90\}$ in the covariate combinations described above (a larger $1-\alpha$ implies a wider region). Figure 3 contains information on age and BMI, while Figure 4 shows gender differences. The FPG variable is plotted on a logarithmic scale. The colored lines indicate the diagnostic thresholds of the American Diabetes Association (ADA): $\mathrm{FPG}\geq 126\penalty 10000\ \mathrm{mg/dL}$ and $\mathrm{HbA1c}\geq 6.5\%$.
We discuss the statistically supported variable by variable. The estimated joint distributions for women and men are similar in shape; however, women exhibit more extreme observation and a higher prevalence of elevated fasting plasma glucose (FPG), especially at higher $\alpha\in(0,1)$ levels. With increasing age, both FPG and HbA1c shift upward, indicating poorer glycemic control at older ages; this effect is most pronounced and consistent near the $0.9$ quantile. A higher body mass index (BMI) is moderately associated with higher glucose levels: people with $\mathrm{BMI}>30\penalty 10000\ \mathrm{kg}/\mathrm{m}^{2}$ show higher values than those with $\mathrm{BMI}\in[20,30]\penalty 10000\ \mathrm{kg}/\mathrm{m}^{2}$. Between $30$ and $40\penalty 10000\ \mathrm{kg}/\mathrm{m}^{2}$, the tolerance regions change little between the quantile levels. In general, glucose control differs between overweight/obesity individuals and those with a BMI in the normal range.
From these results, two implications follow: (i) glycemic control in the U.S. population varies significantly between demographic and anthropometric individual characteristics, with consequences for public-health planning and political interventions; and (ii) diagnostic cutoff points for diabetes based on FPG and HbA1c may warrant adjustment for at least age and possibly BMI, rather than relying on the same rules for all population in line with [8] . From a modeling and analysis perspective, SDRF provides a principled framework to define reference values for specific subpopulations, reclassify individuals based on biomarker profiles, and forecast public–health risks leveraging the full conditional distribution. In contrast, approaches that model only the conditional mean can miss clinically important features—such as quantiles, dispersion, and tail behavior, leading to less informative inform to perform the decision making.
Figure 3: SDRF–estimated tolerance regions (in expectation) for $\log(\mathrm{FPG})$ vs. $\mathrm{HbA1c}$ at quantile levels $\alpha\in\{0.10,0.25,0.50,0.75,0.90\}$. Panels display regions conditional on specific age groups (top two) and BMI ranges (bottom three). Green and gray points indicate observations falling inside and outside the estimated regions, respectively.
Figure 4: SDRF–estimated tolerance regions (in expectation) for $\log(\mathrm{FPG})$ vs. $\mathrm{HbA1c}$ stratified by gender. The contours correspond to $\alpha\in\{0.10,0.25,0.50,0.75,0.90\}$. Green and gray points indicate observations falling inside and outside the estimated regions, respectively.
6 Discussion
We propose a random forest tailored to complex survey designs, in which the bagging step explicitly mimics the sampling design (e.g., strata, primary sampling units, and weights), inspired by recent progress on bootstrap methods for complex designs with rate improvements [95]. To the best of our knowledge, random forests and other tree ensembles have not previously been adapted to complex survey data, filling an important methodological gap. From an estimation and applied data analysis perspective, we introduce what is, to our knowledge, the first general framework for complex designs to estimate conditional distributions for multivariate outcomes. Our approach leverages reproducing-kernel Hilbert spaces (RKHS) and distances between kernel mean embeddings to estimate the full conditional distribution. This representation enables straightforward derivation of population quantities (e.g., quantiles, ratios) and supports downstream tasks–including two-sample testing, selection of distributional variables, and semi-supervised clustering—via a forest-based kernel mean embedding approach [7, 67, 79, 37].
SDRF framework provides a powerful and flexible tool for biomedical studies conducted under complex survey designs, as illustrated by the NHANES case study reported here, studies that are essential for optimizing public-health decisions. Our machine learning algorithm for estimating the conditional distribution of multivariate outcomes represents a substantial modeling advance, enabling investigators to address clinically relevant questions in these cohorts with direct implications for health policy, as demonstrated in our case study illustration.
Survey data and machine learning offer promising synergies for more efficient estimation in finite sample regimes. Building on recent work, we have the potential to improve calibration, reduce overfitting, and improve algorithmic fairness in population-based studies, while mitigating selection bias and preserving practical causal model interpretations-—especially in nationally representative healthcare studies.
In future work, we will continue to investigate the survey-design aspects of our method. In particular, we will study the upward bias that can arise in ratio-type estimators when the denominator (here, the leaf size or subpopulation size) must be estimated because the true leaf sizes are not available from the finite population or census. Extending SDRF for outcome-dependent sampling is another direction.
From a data-analytic perspective, we will develop algorithms to handle non-Euclidean random objects [58] that arise in digital health applications, and we will design methods to quantify uncertainty (for example, via confidence bands) within our random forest frameworks. It is also of applied interest to extend our algorithm to functional data [24], which can be naturally accommodated by using kernel RKHS embeddings to define split rules. There are recent regression methods for functional data in complex survey designs, but they focus only on linear functional response models and the conditional regression mean operator [52, 65].